# Few for Many: Tchebycheff Set Scalarization for Many-Objective Optimization

### v1

@@ 3 Tchebycheff Set Scalarization for Many Objective Optimization
Figure 2: Few Solutions to Address Many Optimization Objectives. (a)-(e): $5$ different solutions to tackle different optimization objectives in a complementary manner. (f): They together successfully handle all $100$ optimization objectives.
@@ 3.1 Small Solution Set for Many-Objective Optimization
$\min_{\boldsymbol{X}_{K}=\{\boldsymbol{x}^{(k)}\}_{k=1}^{K}}\boldsymbol{f}(\boldsymbol{x})=(\min_{\boldsymbol{x}\in\boldsymbol{X}_{K}}f_{1}(\boldsymbol{x}),\min_{\boldsymbol{x}\in\boldsymbol{X}_{K}}f_{2}(\boldsymbol{x}),\cdots,\min_{\boldsymbol{x}\in\boldsymbol{X}_{K}}f_{m}(\boldsymbol{x})),$ (6)
where $\boldsymbol{X}_{K}=\{\boldsymbol{x}^{(k)}\}_{k=1}^{K}$ is a set of $K$ solutions to tackle all $m$ objectives $\{f_{i}(\boldsymbol{x})\}_{i=1}^{m}$. With a large $K\geq m$, we will have a degenerated problem:
$\min_{\boldsymbol{X}_{K}}\boldsymbol{f}(\boldsymbol{x})=(\min_{\boldsymbol{x}^{(1)}\in\mathcal{X}}f_{1}(\boldsymbol{x}),\min_{\boldsymbol{x}^{(2)}\in\mathcal{X}}f_{2}(\boldsymbol{x}),\cdots,\min_{\boldsymbol{x}^{(m)}\in\mathcal{X}}f_{m}(\boldsymbol{x})),$ (7)
where each objective function $f_{i}$ is independently solved by its corresponding solution $\boldsymbol{x}^{(i)}\in\mathcal{X}$ via single objective optimization and the rest $(K-m)$ solutions are redundant. If $K=1$, the set optimization problem (6) will be reduced to the standard multi-objective optimization problem (1).
In this work, we are more interested in the case $1<K\ll m$, which finds a small set of solutions (e.g., $K=5$) to tackle a large number of objectives (e.g., $m\geq 100$) as illustrated in Figure 2. Very recently, a similar setting has been investigated in two concurrent works [14, 31]. Ding et al.[14] study the sum-of-minimum (SoM) optimization problem $\frac{1}{m}\sum_{i=1}^{m}\min\{f_{i}(\boldsymbol{x}^{(1)}),f_{i}(\boldsymbol{x}^{(2)}),\ldots,f_{i}(\boldsymbol{x}^{(K)})\}$ that can be found in many machine learning applications such as mixed linear regression [59, 62]. They generalize the classic k-means++ [2] and Lloyd’s algorithm [41] for clustering to tackle this problem, but do not take multi-objective optimization into consideration. Li et al. [31] propose a novel Many-objective multi-solution Transport (MosT) framework to tackle many-objective optimization. With a bi-level optimization formulation, they adaptively construct a few weighted multi-objective optimization problems that are assigned to different representative regions on the Pareto front. By solving these weighted problems with MGDA, a diverse set of solutions can be obtained to well cover all objectives. In this work, we propose a straightforward and efficient set scalarization approach to explicitly optimize all objectives by a small set of solutions. A detailed experimental comparison with these methods can be found in Section 4.
@@ 3.2 Tchebycheff Set Scalarization
The set optimization formulation (6) is still a multi-objective optimization problem. In non-trivial cases, there is no single small solution set $\boldsymbol{X}_{K}$ with $K<m$ solutions that can optimize all $m$ objective functions $\{f_{i}(\boldsymbol{x})\}_{i=1}^{m}$ at the same time. To tackle this optimization problem, we propose the following Tchebycheff set (TCH-Set) scalarization approach:
$\min_{\boldsymbol{X}_{K}=\{\boldsymbol{x}^{(k)}\}_{k=1}^{K}}g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ $=\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{\boldsymbol{x}\in\boldsymbol{X}_{K}}f_{i}(\boldsymbol{x})-z^{*}_{i})\right\}$
$=\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{1\leq k\leq K}f_{i}(\boldsymbol{x}^{(k)})-z^{*}_{i})\right\},$ (8)
@@
where $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})$ and $\boldsymbol{z}^{*}=(z^{*}_{1},\dots,z^{*}_{m})$ are the preference and ideal point for each objective function. In this way, all objective values $\{f_{i}(\boldsymbol{x})\}_{i=1}^{m}$ among the whole solution set $\boldsymbol{X}_{K}=\{\boldsymbol{x}^{(k)}\}_{k=1}^{K}$ are scalarized into a single function $g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$. A simple uniform vector $\boldsymbol{\lambda}=(\frac{1}{m},\ldots,\frac{1}{m})$ can be used if there is no specific preference among the objectives. By optimizing this TCH-Set scalarization function (8), we want to find an optimal small solution set $\boldsymbol{X}^{*}_{K}$ such that each objective can be well addressed by at least one solution $\boldsymbol{x}^{(k)}\in\boldsymbol{X}_{K}$ with a low worst objective value $\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{1\leq k\leq K}f_{i}(\boldsymbol{x}^{(k)})-z^{*}_{i})\right\}$. When the solution set contains only one single solution (e.g., $K=1$), it will be reduced to the classic single-solution Tchebycheff scalarization (4).
@@
In this work, we assume that no pair of the optimization objectives will share exactly the same optimal solution. Therefore, no solution in the optimal solution set $\boldsymbol{X}^{*}_{K}=\operatorname*{arg\,min}_{\boldsymbol{X}_{k}}g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ will be redundant, which means
$g^{(\text{TCH-Set})}(\boldsymbol{X}^{*}_{K}|\boldsymbol{\lambda})<g^{(\text{TCH-Set})}(\boldsymbol{X}^{*}_{K}\setminus\{\boldsymbol{x}^{(k)}\}|\boldsymbol{\lambda})$ (9)
holds for any $1\leq k\leq K\leq m$ with a all positive preference $\boldsymbol{\lambda}$. In addition, the optimal solution set also enjoys a good theoretical property for multi-objective optimization:
Theorem 1 (Pareto Optimality for Tchebycheff Set Scalarization).
@@
All solutions in the optimal solution set $\boldsymbol{X}^{*}_{K}$ for the Tchebycheff set scalarization optimization problem (8) are weakly Pareto optimal of the original multi-objective optimization problem (1). In addition, if the optimal set $\boldsymbol{X}^{*}_{K}$ is unique, all solutions in $\boldsymbol{X}^{*}_{K}$ are Pareto optimal.
@@
This theorem can be proved by contradiction based on Definition 2 for (weakly) Pareto optimality and the form of Tchebycheff set Scalarization (8). A detailed proof is provided in Appendix A.1. ∎
In other words, a small set of (weakly) Pareto solutions can be found by solving the TCH-Set scalarization optimization problem (8), which complementarily addresses all objectives to achieve a low worst performance $\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{1\leq k\leq K}f_{i}(\boldsymbol{x}^{(k)})-z^{*}_{i})\right\}$. However, the non-smoothness of TCH-Set scalarization might hinder its practical usage for efficient optimization.
@@ 3.3 Smooth Tchebycheff Set Scalarization
The Tchebycheff set scalarization formulation (8) involves a $\max$ and a $\min$ operator, which leads to its non-smoothness even when all objective functions $\{f_{i}(\boldsymbol{x})\}_{i=1}^{m}$ are smooth. In other words, the Tchebycheff set scalarization $g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ is not differentiable and will suffer from a slow convergence rate by subgradient descent. To address this issue, we leverage the smooth optimization approach [48, 4, 9] to propose a smooth Tchebycheff set scalarization for multi-objective optimization.
According to [4], for the maximization among all objectives $\max_{1\leq i\leq m}\{f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\ldots,f_{m}(\boldsymbol{x})\}$, we have the smooth maximization function:
$\operatorname*{smax}_{1\leq i\leq m}\{f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\ldots,f_{m}(\boldsymbol{x})\}=\mu\log\left(\sum_{i=1}^{m}e^{\frac{f_{i}(\boldsymbol{x})}{\mu}}\right),$ (10)
where $\mu$ is a smooth parameter. Similarly, for the minimization among different solutions for the $i$-th objective $\min_{1\leq k\leq K}\{f_{i}(\boldsymbol{x}^{(1)}),f_{i}(\boldsymbol{x}^{(2)}),\ldots,f_{i}(\boldsymbol{x}^{(K)})\}$, we have the smooth minimization function:
$\operatorname*{smin}_{1\leq k\leq K}\{f_{i}(\boldsymbol{x}^{(1)}),f_{i}(\boldsymbol{x}^{(2)}),\ldots,f_{i}(\boldsymbol{x}^{(K)})\}=-\mu_{i}\log\left(\sum_{k=1}^{K}e^{-\frac{f_{i}(\boldsymbol{x}^{(k)})}{\mu_{i}}}\right),$ (11)
@@
$\min_{\boldsymbol{X}_{K}=\{\boldsymbol{x}^{(k)}\}_{k=1}^{K}}g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ $=\operatorname*{smax}_{1\leq i\leq m}\left\{\lambda_{i}\operatorname*{smin}\limits_{1\leq k\leq K}f_{i}(\boldsymbol{x}^{(k)})-z^{*}_{i}\right\}$
$=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}\left(\operatorname*{smin}\limits_{1\leq k\leq K}f_{i}(\boldsymbol{x}^{(k)})-z^{*}_{i}\right)}{\mu}}\right)$
$=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}\left(-\mu_{i}\log\left(\sum_{k=1}^{K}e^{-{f_{i}(\boldsymbol{x}^{(k)})}/{\mu_{i}}}\right)-z^{*}_{i}\right)}{\mu}}\right).$ (12)
@@
$\min_{\boldsymbol{X}_{K}=\{\boldsymbol{x}^{(k)}\}_{k=1}^{K}}g_{\mu}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{\lambda_{i}\left(-\log\left(\sum_{k=1}^{K}e^{-{f_{i}(\boldsymbol{x}^{(k)})}/{\mu}}\right)-z^{*}_{i}\right)}\right).$ (13)
@@
All solutions in the optimal solution set $\boldsymbol{X}^{*}_{K}$ for the smooth Tchebycheff set scalarization problem (12) are weakly Pareto optimal of the original multi-objective optimization problem (1). In addition, the solutions are Pareto optimal if either
@@
    the optimal solution set $\boldsymbol{X}^{*}_{K}$ is unique, or
@@
The smooth Tchebycheff set (STCH-Set) scalarization $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ is a uniform smooth approximation of the Tchebycheff set (TCH-Set) scalarization $g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$, and we have:
$\lim_{\mu\downarrow 0,\mu_{i}\downarrow 0\ \forall i}g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})=g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ (14)
for any valid set $\boldsymbol{X}_{k}\subset\mathcal{X}$.
@@
This theorem can be proved by deriving the upper and lower bounds of the TCH-Set scalarization $g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ with respect to the smooth $\operatorname*{smax}$ and $\operatorname*{smin}$ operators in the STCH-Set scalarization $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$. A detailed proof is provided in Appendix A.3. ∎
@@
1:  Input: Preference $\boldsymbol{\lambda}$, Step Size $\{\eta_{t}\}_{t=0}^{T}$, Initial $\boldsymbol{X}^{0}_{K}$
@@
3:     $\boldsymbol{X}^{t}_{K}=\boldsymbol{X}^{t-1}_{K}-\eta_{t}\nabla g^{(\text{STCH-Set})}(\boldsymbol{X}^{t-1}_{K}|\boldsymbol{\lambda})$
@@
5:  Output: Final Solution $\boldsymbol{X}^{T}_{K}$
According to these theorems, for any smooth parameters $\mu$ and $\{\mu\}_{i=1}^{m}$, all solutions in the optimal set $\boldsymbol{X}^{*}_{K}$ for STCH-Set scalarization (12) are also (weakly) Pareto optimal. In addition, with small smooth parameters $\mu\downarrow 0,\mu_{i}\downarrow 0\ \forall i$, the value of $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ will be close to its non-smooth counterpart $g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ for all valid solution sets $\boldsymbol{X}_{K}$, which is also the case for the optimal set $\boldsymbol{X}^{*}_{K}=\operatorname*{arg\,min}_{\boldsymbol{X}_{K}}g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$. Therefore, it is reasonable to find an approximate optimal solution set for the original $g^{(\text{TCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ by optimizing its smooth counterpart $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ with small smooth parameters $\mu$ and $\{\mu_{i}\}_{i=1}^{m}$.
The STCH-Set scalarization function $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$ can be efficiently optimized by any gradient-based optimization method. A simple gradient descent algorithm for STCH-Set is shown in Algorithm 1. However, when the objective functions $\{f_{i}(\boldsymbol{x})\}_{i=1}^{m}$ are highly non-convex, it could be very hard to find the global optimal set $\boldsymbol{X}^{*}_{K}$ for $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\boldsymbol{X}_{K}|\boldsymbol{\lambda})$. In this case, we have the Pareto stationarity guarantee for the gradient-based method:
@@
If there exists a solution set ^$\hat{}\boldsymbol{X}_{K}$ such that ^$\nabla_{\hat{}\boldsymbol{x}^{(k)}}g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\hat{}\boldsymbol{X}_{K}|\boldsymbol{\lambda})=\bm{0}$ for all ^$\hat{}\boldsymbol{x}^{(k)}\in\hat{}\boldsymbol{X}_{K}$, then all solutions in ^$\hat{}\boldsymbol{X}_{K}$ are Pareto stationary solutions of the original multi-objective optimization problem (1).

### v3

@@ 3 Tchebycheff Set Scalarization for Many Objective Optimization
Figure 2: Few Solutions to Address Many Optimization Objectives. (a)-(e): 55 different solutions to tackle different optimization objectives in a complementary manner. (f): They together successfully handle all $100$ optimization objectives.
@@ 3.1 Small Solution Set for Many-Objective Optimization
$\min_{\bm{X}_{K}=\{\bm{x}^{(k)}\}_{k=1}^{K}}\bm{f}(\bm{x})=(\min_{\bm{x}\in\bm{X}_{K}}f_{1}(\bm{x}),\min_{\bm{x}\in\bm{X}_{K}}f_{2}(\bm{x}),\cdots,\min_{\bm{x}\in\bm{X}_{K}}f_{m}(\bm{x})),$ (6)
where $\bm{X}_{K}=\{\bm{x}^{(k)}\}_{k=1}^{K}$ is a set of $K$ solutions to tackle all $m$ objectives $\{f_{i}(\bm{x})\}_{i=1}^{m}$. With a large $K\geq m$, we will have a degenerated problem:
$\min_{\bm{X}_{K}}\bm{f}(\bm{x})=(\min_{\bm{x}^{(1)}\in\mathcal{X}}f_{1}(\bm{x}),\min_{\bm{x}^{(2)}\in\mathcal{X}}f_{2}(\bm{x}),\cdots,\min_{\bm{x}^{(m)}\in\mathcal{X}}f_{m}(\bm{x})),$ (7)
where each objective function $f_{i}$ is independently solved by its corresponding solution $\bm{x}^{(i)}\in\mathcal{X}$ via single objective optimization and the rest (Km)(K-m) solutions are redundant. If $K=1$, the set optimization problem (6) will be reduced to the standard multi-objective optimization problem (1).
In this work, we are more interested in the case $1<K\ll m$, which finds a small set of solutions (e.g., $K=5$) to tackle a large number of objectives (e.g., $m\geq 100$) as illustrated in Figure 2. In the ideal case, if the ground truth optimal objective group assignment is already known (e.g., which objectives should be optimized together by the same solution), it is straightforward to directly find an optimal solution for each group of objectives. However, for a general optimization problem, the ground truth objective group assignment is usually unknown in most cases, and finding the optimal assignment could be very difficult.
Very recently, a similar setting has been investigated in two concurrent works (Ding et al., 2024; Li et al., 2024). Ding et al. (2024) study the sum-of-minimum (SoM) optimization problem $\frac{1}{m}\sum_{i=1}^{m}\min\{f_{i}(\bm{x}^{(1)}),f_{i}(\bm{x}^{(2)}),\ldots,f_{i}(\bm{x}^{(K)})\}$ that can be found in many machine learning applications such as mixed linear regression (Yi et al., 2014; Zhong et al., 2016). They generalize the classic k-means++ (Arthur, 2007) and Lloyd’s algorithm (Lloyd, 1982) for clustering to tackle this problem, but do not take multi-objective optimization into consideration. Li et al. (2024) propose a novel Many-objective multi-solution Transport (MosT) framework to tackle many-objective optimization. With a bi-level optimization formulation, they adaptively construct a few weighted multi-objective optimization problems that are assigned to different representative regions on the Pareto front. By solving these weighted problems with MGDA, a diverse set of solutions can be obtained to well cover all objectives. In this work, we propose a straightforward and efficient set scalarization approach to explicitly optimize all objectives by a small set of solutions. A detailed experimental comparison with these methods can be found in Section 4.
@@ 3.2 Tchebycheff Set Scalarization
The set optimization formulation (6) is still a multi-objective optimization problem. In non-trivial cases, there is no single small solution set $\bm{X}_{K}$ with $K<m$ solutions that can optimize all $m$ objective functions $\{f_{i}(\bm{x})\}_{i=1}^{m}$ at the same time. To tackle this optimization problem, we propose the following Tchebycheff set (TCH-Set) scalarization approach:
$\min_{\bm{X}_{K}=\{\bm{x}^{(k)}\}_{k=1}^{K}}g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ $=\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{\bm{x}\in\bm{X}_{K}}f_{i}(\bm{x})-z^{*}_{i})\right\}$
$=\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{1\leq k\leq K}f_{i}(\bm{x}^{(k)})-z^{*}_{i})\right\},$ (8)
@@
where $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{m})$ and $\bm{z}^{*}=(z^{*}_{1},\dots,z^{*}_{m})$ are the preference and ideal point for each objective function. In this way, all objective values $\{f_{i}(\bm{x})\}_{i=1}^{m}$ among the whole solution set $\bm{X}_{K}=\{\bm{x}^{(k)}\}_{k=1}^{K}$ are scalarized into a single function $g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$. In this work, a simple uniform vector $\bm{\lambda}=(\frac{1}{m},\ldots,\frac{1}{m})$ is used in all experiments without any specific preference among the objectives. A discussion on the effect of different preferences can be found in Appendix D.2.
By optimizing this TCH-Set scalarization function (8), we want to find an optimal small solution set $\bm{X}^{*}_{K}$ such that each objective can be well addressed by at least one solution $\bm{x}^{(k)}\in\bm{X}_{K}$ with a low worst objective value $\max_{1\leq i\leq m}\left\{\lambda_{i}(\min_{1\leq k\leq K}f_{i}(\bm{x}^{(k)})-z^{*}_{i})\right\}$. When the solution set contains only one single solution (e.g., $K=1$), it will be reduced to the classic single-solution Tchebycheff scalarization (4). To avoid degenerated cases and focus on the key few-for-many setting, we make the following two assumptions in this work:
Assumption 1 (No Redundant Solution).
@@
We assume that no solution in the optimal solution set $X^{*}_{K}=\mathtt{argmin}_{X_{k}}g^{(\text{TCH-Set})}(X_{K}|\lambda)$ will be redundant, which means
$g^{(\text{TCH-Set})}(X^{*}_{K}|\lambda)<g^{(\text{TCH-Set})}(X^{*}_{K}\setminus\{x^{(k)}\}|\lambda)$ (9)
holds for any $1\leq k\leq K\leq m$ with $\lambda>\bm{0}$.
Assumption 2 (All Positive Preference).
In the few-for-many setting, we assume that the preferences should be positive for all objectives such that $\{\sum_{i=1}^{m}\lambda_{i}=1,\lambda_{i}>0\ \forall i\}$.
These assumptions are reasonable in practice, especially for the few-for-many setting (e.g., $K\ll m$) considered in this work. For a general problem, it is extremely rare that a small number of solutions (e.g., $4$) can exactly solve a large number of objectives (e.g., $m=1,000$). From the viewpoint of clustering (Ding et al., 2024), it is analogous to the case that all the $1,000$ data points under consideration are exactly located in $4$ different locations. If we do meet this extreme case with $p$ redundant solutions, we can simply reduce the number of solutions to make it a (Kp)(K-p)-solutions-for-$m$-objectives problem. Similarly, if some of the preferences (e.g., $q$) are 0, we can also simply reduce the original $m$-objective problem into a (mq)(m-q)-objective problem that satisfies the positive preference assumption.
From the viewpoint of multi-objective optimization, we have the following guarantee for the optimal solution set of Tchebycheff set scalarization:
Theorem 1 (Existence of Pareto Optimal Solution for Tchebycheff Set Scalarization).
@@
There exists an optimal solution set $\bar{}\bm{X}^{*}_{K}$ for the Tchebycheff set scalarization optimization problem (8) such that all solutions in $\bar{}\bm{X}^{*}_{K}$ are Pareto optimal of the original multi-objective optimization problem (1). In addition, if the optimal set $\bm{X}^{*}_{K}$ is unique, all solutions in $\bm{X}^{*}_{K}$ are Pareto optimal.
@@
This theorem can be proved by construction and contradiction based on Definition 2 for (weakly) Pareto optimality and the form of Tchebycheff set Scalarization (8). A detailed proof is provided in Appendix A.1. ∎
It should be emphasized that, without the strong unique optimal solution set assumption, we only have a weak existence guarantee for the Pareto optimality. For an optimal solution set for the Tchebycheff set scalarization (8), it is possible that many of the solutions are not even weakly Pareto optimal for the original multi-objective optimization problem (1). This finding could be a bit surprising for multi-objective optimization, and we provide a detailed discussion in Appendix A.5. In addition to the Pareto optimality guarantee, the non-smoothness of TCH-Set scalarization might also hinder its practical usage for efficient optimization. To address these crucial issues, we further develop a smooth Tchebycheff set scalarization with good optimization property and promising Pareto optimality guarantee in the next subsection.
@@ 3.3 Smooth Tchebycheff Set Scalarization
The Tchebycheff set scalarization formulation (8) involves a $\max$ and a $\min$ operator, which leads to its non-smoothness even when all objective functions $\{f_{i}(\bm{x})\}_{i=1}^{m}$ are smooth. In other words, the Tchebycheff set scalarization $g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ is not differentiable and will suffer from a slow convergence rate by subgradient descent. To address this issue, we leverage the smooth optimization approach (Nesterov, 2005; Beck and Teboulle, 2012; Chen, 2012) to propose a smooth Tchebycheff set scalarization for multi-objective optimization.
According to Beck and Teboulle (2012), for the maximization function among all objectives $\max_{1\leq i\leq m}\{f_{1}(\bm{x}),f_{2}(\bm{x}),\ldots,f_{m}(\bm{x})\}$, we have the smooth maximization function:
$\operatorname*{smax}_{1\leq i\leq m}\{f_{1}(\bm{x}),f_{2}(\bm{x}),\ldots,f_{m}(\bm{x})\}=\mu\log\left(\sum_{i=1}^{m}e^{\frac{f_{i}(\bm{x})}{\mu}}\right),$ (10)
where $\mu$ is a smooth parameter. Similarly, for the minimization among different solutions for the $i$-th objective $\min_{1\leq k\leq K}\{f_{i}(\bm{x}^{(1)}),f_{i}(\bm{x}^{(2)}),\ldots,f_{i}(\bm{x}^{(K)})\}$, we have the smooth minimization function:
$\operatorname*{smin}_{1\leq k\leq K}\{f_{i}(\bm{x}^{(1)}),f_{i}(\bm{x}^{(2)}),\ldots,f_{i}(\bm{x}^{(K)})\}=-\mu_{i}\log\left(\sum_{k=1}^{K}e^{-\frac{f_{i}(\bm{x}^{(k)})}{\mu_{i}}}\right),$ (11)
@@
$\min_{\bm{X}_{K}=\{\bm{x}^{(k)}\}_{k=1}^{K}}g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ $=\operatorname*{smax}_{1\leq i\leq m}\left\{\lambda_{i}\operatorname*{smin}\limits_{1\leq k\leq K}f_{i}(\bm{x}^{(k)})-z^{*}_{i}\right\}$
$=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}\left(\operatorname*{smin}\limits_{1\leq k\leq K}f_{i}(\bm{x}^{(k)})-z^{*}_{i}\right)}{\mu}}\right)$
$=\mu\log\left(\sum_{i=1}^{m}e^{\frac{\lambda_{i}\left(-\mu_{i}\log\left(\sum_{k=1}^{K}e^{-{f_{i}(\bm{x}^{(k)})}/{\mu_{i}}}\right)-z^{*}_{i}\right)}{\mu}}\right).$ (12)
@@
$\min_{\bm{X}_{K}=\{\bm{x}^{(k)}\}_{k=1}^{K}}g_{\mu}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})=\mu\log\left(\sum_{i=1}^{m}e^{\lambda_{i}\left(-\log\left(\sum_{k=1}^{K}e^{-{f_{i}(\bm{x}^{(k)})}/{\mu}}\right)-z^{*}_{i}\right)}\right).$ (13)
@@
All solutions in the optimal solution set $\bm{X}^{*}_{K}$ for the smooth Tchebycheff set scalarization problem (12) are weakly Pareto optimal of the original multi-objective optimization problem (1). In addition, the solutions are Pareto optimal if either
@@
    the optimal solution set $\bm{X}^{*}_{K}$ is unique, or
@@
The smooth Tchebycheff set (STCH-Set) scalarization $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ is a uniform smooth approximation of the Tchebycheff set (TCH-Set) scalarization $g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$, and we have:
$\lim_{\mu\downarrow 0,\mu_{i}\downarrow 0\ \forall i}g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})=g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ (14)
for any valid set $\bm{X}_{k}\subset\mathcal{X}$.
@@
This theorem can be proved by deriving the upper and lower bounds of the TCH-Set scalarization $g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ with respect to the smooth $\operatorname*{smax}$ and $\operatorname*{smin}$ operators in the STCH-Set scalarization $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$. A detailed proof is provided in Appendix A.3. ∎
According to these theorems, for any smooth parameters $\mu$ and $\{\mu\}_{i=1}^{m}$, all solutions in the optimal set $\bm{X}^{*}_{K}$ for STCH-Set scalarization (12) are also (weakly) Pareto optimal. In addition, with small smooth parameters $\mu\downarrow 0,\mu_{i}\downarrow 0\ \forall i$, the value of $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ will be close to its non-smooth counterpart $g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ for all valid solution sets $\bm{X}_{K}$, which is also the case for the optimal set $\bm{X}^{*}_{K}=\operatorname*{arg\,min}_{\bm{X}_{K}}g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$. Therefore, it is reasonable to find an approximate optimal solution set for the original $g^{(\text{TCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ by optimizing its smooth counterpart $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ with small smooth parameters $\mu$ and $\{\mu_{i}\}_{i=1}^{m}$.
@@
1:Input: Preference $\bm{\lambda}$, Step Size $\{\eta_{t}\}_{t=0}^{T}$, Initial $\bm{X}^{0}_{K}$
@@
3:  $\bm{X}^{t}_{K}=\bm{X}^{t-1}_{K}-\eta_{t}\nabla g^{(\text{STCH-Set})}(\bm{X}^{t-1}_{K}|\bm{\lambda})$
@@
5:Output: Final Solution $\bm{X}^{T}_{K}$
The STCH-Set scalarization function $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$ can be efficiently optimized by any gradient-based optimization method, where we treat all solutions as a single solution matrix. A simple gradient descent algorithm for STCH-Set is shown in Algorithm 1. However, when the objective functions $\{f_{i}(\bm{x})\}_{i=1}^{m}$ are highly non-convex, it could be very hard to find the global optimal set $\bm{X}^{*}_{K}$ for $g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\bm{X}_{K}|\bm{\lambda})$. In this case, we have the Pareto stationarity guarantee for the gradient-based method:
@@
If there exists a solution set $\hat{}\bm{X}_{K}$ such that $\nabla_{\hat{}\bm{x}^{(k)}}g_{\mu,\{\mu_{i}\}_{i=1}^{m}}^{(\text{STCH-Set})}(\hat{}\bm{X}_{K}|\bm{\lambda})=\bm{0}$ for all $\hat{}\bm{x}^{(k)}\in\hat{}\bm{X}_{K}$, then all solutions in $\hat{}\bm{X}_{K}$ are Pareto stationary solutions of the original multi-objective optimization problem (1).